Company A purchased a patent for $60,000. The patent has a legal life of 5 year, and it will generate benefit for 20 years, but company A plans to sell it at the end of 3rd year. What is the book value of the patent at the end of year 3?
Answer #1

Amortization per year = (60000-0)/5 = $12000

Book value of the patent at the end of year 3 = $60000 - 12000*3 = $24000